OpenStudy (anonymous):

1- What is the formula for the sum of the interior angles of a polygon? 2- What is the interior angle sum of an octagon (8 Sides) 3-What is the measure of one interior angle of a hexagon (6 Sides) 4- What is the measure of one exterior angle of a regular 10-gon

OpenStudy (anonymous):

\[\huge 180(n-2) \]

OpenStudy (jdoe0001):

\(\bf \textit{sum of interior angles}=180(n-2)\qquad n=sides\)

OpenStudy (jdoe0001):

2- to get it for an octagon, well, 8 sides, n = 8 3- 6 sides, n = 6 4- 10 sides, n = 10

OpenStudy (anonymous):

So 180*6 for 2

OpenStudy (jdoe0001):

yeap 180 * (8-2) 180 * 6

OpenStudy (anonymous):

So 1080 is the answer